# Unified Locomotion Transformer with Simultaneous Sim-to-Real Transfer for Quadrupeds

### v1

@@ I Introduction
Driven by the advancement of deep reinforcement learning (RL), quadruped robots have drawn great attention, due to their capability of traversing across complex terrains [1, 2, 3, 4]. Traditional robotics controllers rely on dedicated models and heuristics, which require extensive prior knowledge and can struggle to adapt to dynamic environments and unpredictable situations [3, 4]. Recently, a new paradigm is introduced to push quadrupeds’ limit to handle complex tasks in challenging environments [5, 6, 7, 8]: a controller is learned from simulations that contain various environmental and physical factors, and then transferred to the physical robot through RL and various knowledge transfer techniques [9, 3, 4, 10, 11] to overcome the sim-to-real gap using privilege information.
One of the commonly used knowledge distillation methods is the teacher-student framework [3]. A teacher policy is first trained through RL, and privilege information about the environment is provided for learning the optimal locomotion strategies efficiently. As additional information, including ground-truth dynamics and terrain profiles, is often inaccessible in the real world, a student policy is subsequently trained to make the model deployable. This is typically achieved through supervised knowledge distillation, either in an offline [3] or online [4] fashion by creating a data set with new trajectories and associated action labels of the teacher created through algorithms such as Data Aggregator (DAgger) [12].
@@
Motivated by the above limitation, the objective of this paper is to simplify the knowledge distillation process with a unified architecture to keep teacher and student policies in a single network for simultaneous optimization. Our inspiration comes from the exceptional capabilities of transformers in multimodal modeling of temporal and sensory information and context understanding [17], which can well support the policy optimization [14]. Thus, we can introduce the privilege information into the observation as another modality to form a unified framework for single-phase optimization of teacher and student policies, achieving zero-shot sim-real transfer.
To this end, we propose Unified Locomotion Transformer (ULT), a new unified framework for end-to-end quadruped locomotion. It is based on the standard transformer architecture with casual masking to pack teacher and student policies in a single network. These policies are optimized jointly with reinforcement learning, next state-action prediction, and action imitation, all in just one training phase, to overcome the sim-to-real gap and achieve zero-shot deployment. In this way, we eliminate the need of dedicated design and training of a teacher network, while the privilege information can still efficiently guide the student policy with proprioception observation during the exploration to generate more diverse trajectories to improve the overall generalization and robustness.
@@ III-A Terrain and Curriculum
To ensure that the policy is robust against different indoor and outdoor environments, we adopt the terrain curriculum from [25] with smooth slope, rough slope, stairs up, stairs down and discrete obstacle terrains. Each type of terrain has 10 levels with incremental difficulty and an overall proportion of $[0.1,0.1,0.35,0.25,0.2]$, respectively. The linear velocity return is tracked across each trajectory’s life cycle. The level is considered solved when an agent reaches 80% of the maximum achievable tracking reward and progresses to the next level. If any agent fails to reach 25% of the maximum reward, it will regress to a lower level for more exploration.
@@ III-B Domain Randomization
Following  [25, 14], we apply Domain Randomization (DR) on key dynamics parameters to enhance the robustness of the policy. To simulate the actual user commands, we sample the linear commands in longitudinal and lateral direction separately with a uniform distribution in $[-0.5,0.5]$ m/s. For the angular command, we first sample the desired heading of the robot and cap the resulted angular velocity command at 0.5 rad/s. The commands are resampled every 10 seconds. Table I lists the key parameters of DR used in the simulation.
@@ III-C Observations and Actions
Privilege Information. To achieve an optimal locomotion policy with the hidden information about the environment, related privilege data is extracted from simulation to form the privilege observation $e_{t}$ for the teacher policy to utilize. The privilege information contains randomized dynamics parameters $d_{t}$ sampled from Sec. III-B, ground truth robot states $s_{t}$ including base velocity, orientation, and precise surrounding height map $m_{t}$. Although such information is often inaccessible in the real world, it helps the teacher policy reconstruct the states and improve the learning efficiency [3, 4].
Proprioceptive Observation. In order to conduct knowledge transfer for a deployable agent, the student policy only relies on onboard sensors to provide observations. Typically, quadrupedal robots are equipped with joint encoders, IMU, and foot contact sensors, which can provide information on joint position $q\in\mathbb{R}^{12}$, joint velocity $\dot{q}\in\mathbb{R}^{12}$, angular velocity $\omega\in\mathbb{R}^{3}$, gravity vector $g\in\mathbb{R}^{3}$ and binary foot contact $c\in\mathbb{R}^{4}$. In order to follow the user command, the agent also needs access to the randomly sampled $\mathrm{cmd}=[v_{x},v_{y},\omega_{z}]\in\mathbb{R}^{3}$ to form the observation of each step $o_{t}=[q,\dot{q},\omega,g,c,\mathrm{cmd}]\in\mathbb{R}^{37}$. To provide the state transition and temporal information, the actions of the previous step $a_{t-1}\in\mathbb{R}^{12}$ are added with a list of historical information $\mathcal{T}=[a_{0},o_{1},a_{1},o_{2},\cdots,a_{t-1},o_{t}]$. We use a rolling window of $t=15$, resulting in a full observation in the space of $\mathbb{R}^{49\times 15}$.
Actions. As typical RL locomotion policies perform inference at the frequency of 50-100 Hz, both the teacher and student policies output the desired joint position $a_{t}$, which is passed to a PD controller running at a much higher frequency for a smooth torque output:
@@ IV Methodology
However, such information is often inaccessible in the real world, and the student policy can only rely on the noisy proprioceptive observation $o_{t}\in\Omega$ during deployment. Essentially, knowledge transfer aims to find the observation probabilities $\mathcal{O}:\mathcal{S}\rightarrow\Omega$ such that the student policy can estimate its current status for decision making, either with latent space estimation [4] or direct action imitation [3, 10, 11].
@@ IV-C Direct Sim-to-Real Deployment
The transformer architecture gives the flexibility of variable input length. During training, causal masking ensures that future information will not be seen by previous tokens and the privilege information will never be leaked to proprioceptive observation. Thus, when deployed in the real world, we can directly remove the last privileged environmental token $h_{e}$ safely from the input sequence to generate the student action $\hat{a}_{i}$ directly with only the historical information $\mathcal{T}$.
@@
  Terrain Metric ULT (Ours) Supervised Transfer Joint Transfer CENet PPO
@@ V-B Comparison with Baselines
  • PPO. We use vanilla PPO [31] to train ULT solely on proprioceptive observation and RL loss, which is equivalent to an action mixer ratio of $\alpha=0$, and with the teacher head and other loss modules disabled.
The comparison results are summarized in Table IV. Although ULT only uses the same number of trajectories as Oracle training, both the teacher and student policies achieve the similar performance level of Oracle performance and outperform other baseline models, which require many more trajectories to be generated for most cases. This shows the high efficiency of our proposed framework. With the increased difficulty of using omnidirectional control on all types of terrain, single-stage offline supervised transfer cannot efficiently capture the dynamics of the environment with just good trajectories, and the rough terrains make it hard for the student to survive. Although online supervised transfer and two-stage transfer significantly improve performance, ULT still outperforms supervised transfer with fewer trajectories used. Joint transfer require manual tuning of the joint ratio and struggles to handle the increased environmental challenges with similar performance to vanilla PPO.
@@ VI Conclusion
Although we have simplified the training process as one single phase without training a dedicated teacher policy, we still need to retrain the model when new task requirement is raised. It is appealing to explore the continual learning and generalization capability through the power of large language models (LLMs) in handling wide variety of information with multimodality as future work.

### v2

@@ I Introduction
Driven by the advancement of deep reinforcement learning (RL), quadruped robots have drawn great attention, due to their capability of traversing across complex terrains [1, 2, 3, 4]. Traditional robotics controllers rely on dedicated models and heuristics, which require extensive prior knowledge and can struggle to adapt to dynamic environments and unpredictable situations. Recently, a new paradigm is introduced to push quadrupeds’ limit to handle complex tasks in challenging environments [3, 4, 5, 6, 7, 8]: a controller is learned from simulations that contain various environmental and physical factors, and then transferred to the physical robot through RL and various knowledge transfer techniques [9, 3, 4, 10, 11] to overcome the sim-to-real gap using privilege information.
One of the commonly used knowledge distillation methods is the teacher-student framework [3]. A teacher policy is first trained through RL, and privilege information about the environment is provided for learning the optimal locomotion strategies efficiently. As additional information, including ground-truth dynamics and terrain profiles, is often inaccessible in the real world, a student policy is subsequently trained to make the model deployable. This is typically achieved through supervised knowledge distillation, either offline [3] or online [4] fashion by creating a data set with new trajectories and associated teacher action labels through algorithms such as the Data Aggregator (DAgger) [12].
@@
Motivated by the above limitation, the objective of this paper is to simplify the knowledge distillation process with a unified architecture to keep teacher and student policies in a single network for simultaneous optimization. Our inspiration comes from the exceptional capabilities of transformers in multimodal modeling of temporal and sensory information and context understanding [17], which can well support policy optimization [14]. Thus, we can introduce the privilege information into the observation as another modality to form a unified framework for single-phase optimization of teacher and student policies, achieving zero-shot sim-real transfer.
To this end, we propose Unified Locomotion Transformer (ULT), a new unified framework for end-to-end quadruped locomotion. It is based on the standard transformer architecture with casual masking to pack teacher and student policies in a single network. These policies are optimized jointly with reinforcement learning, next state-action prediction, and action imitation, all in just one training phase, to overcome the sim-to-real gap and achieve zero-shot deployment. In this way, we eliminate the need for dedicated design and training of a teacher network, while the privilege information can still efficiently guide the student policy during the exploration to generate more diverse trajectories to improve the overall generalization and robustness.
@@ III-A Terrain and Curriculum
To ensure that the policy is robust against different indoor and outdoor environments, we adopt the terrain curriculum from [25] with smooth slope, rough slope, stairs up, stairs down and discrete obstacle terrains. Each type of terrain has 10 levels with incremental difficulty and an overall proportion of [0.1,0.1,0.35,0.25,0.2][0.1,0.1,0.35,0.25,0.2][ 0.1 , 0.1 , 0.35 , 0.25 , 0.2 ], respectively. The linear velocity return is tracked across each trajectory’s life cycle. The level is considered solved when an agent reaches 80% of the maximum achievable tracking reward and progresses to the next level. If any agent fails to reach 25% of the maximum reward, it will regress to a lower level.
@@ III-B Domain Randomization
Following  [25, 14], we apply Domain Randomization (DR) on key dynamics parameters to enhance the robustness of the policy. To simulate the actual user commands, we sample the linear commands in longitudinal and lateral direction separately with a uniform distribution in [0.5,0.5][-0.5,0.5][ - 0.5 , 0.5 ] m/s. For the angular command, we first sample the desired heading of the robot and cap the resulted angular velocity command at 0.5 rad/s. The commands are resampled every 10 seconds. Table I lists the key parameters of DR used in the simulation environment.
@@ III-C Observations and Actions
Privilege Information. To achieve an optimal locomotion policy with the hidden information about the environment, related privilege data is extracted from simulation to form the privilege observation $e_{t}$ for the teacher policy to utilize. The privilege information contains randomized dynamics parameters $d_{t}$ sampled from Sec. III-B, ground truth robot states $s_{t}$ including base velocity, orientation, and precise surrounding height map $m_{t}$. Although such information is often inaccessible in the real world, it helps the teacher policy reconstruct states and improve learning efficiency [3, 4].
Proprioceptive Observation. In order to conduct knowledge transfer for a deployable agent, the student policy only relies on onboard sensors to provide observations. Typically, quadrupedal robots are equipped with multiple sensors, including joint encoders, IMU, and foot contact sensors, which can provide information on joint position $q\in\mathbb{R}^{12}$, joint velocity $\dot{q}\in\mathbb{R}^{12}$, angular velocity $\omega\in\mathbb{R}^{3}$, gravity vector $g\in\mathbb{R}^{3}$ and binary foot contact $c\in\mathbb{R}^{4}$. In order to follow the user command, the agent also needs access to the randomly sampled $\mathrm{cmd}=[v_{x},v_{y},\omega_{z}]\in\mathbb{R}^{3}$ to form the observation of each step $o_{t}=[q,\dot{q},\omega,g,c,\mathrm{cmd}]\in\mathbb{R}^{37}$. To provide the state transition and temporal information, the actions of the previous step $a_{t-1}\in\mathbb{R}^{12}$ are added with a list of historical information $\mathcal{T}=[a_{0},o_{1},a_{1},o_{2},\cdots,a_{t-1},o_{t}]$. We use a rolling window of $t=15$, resulting in a full observation in the space of $\mathbb{R}^{49\times 15}$.
Actions. As typical RL locomotion policies perform inference at the frequency of 50-100 Hz, both the teacher and student policies output the desired joint position $q=a_{t}$, which is passed to a PD controller running at a much higher frequency for a smooth torque output:
@@ IV Methodology
However, such information is often inaccessible in the real world, and the student policy can only rely on noisy proprioceptive observation $o_{t}\in\Omega$ during deployment. Essentially, knowledge transfer aims to find the observation probabilities $\mathcal{O}:\mathcal{S}\rightarrow\Omega$ such that the student policy can estimate its current status for decision making, either with latent space estimation [4] or direct action imitation [3, 10, 11].
@@ IV-C Direct Sim-to-Real Deployment
The transformer architecture gives flexibility in variable input length. During training, causal masking ensures that future information will not be seen by previous tokens and the privilege information will never be leaked to proprioceptive observation. Thus, when deployed in the real world, we can directly remove the last privileged environmental token $h_{e}$ safely from the input sequence to generate the student action $\hat{a}_{i}$ directly with only the historical information $\mathcal{T}$.
@@
  Terrain Metric ULT (Ours) Supervised Transfer
Joint
Transfer
CENet PPO
@@ V-B Comparison with Baselines
  • PPO. We use vanilla PPO [31] to train ULT solely on proprioceptive observation and RL loss, which is equivalent to an action mixer ratio of $\alpha=0$, and with the teacher head and other loss modules disabled, resulting in a student-only training.
The comparison results are summarized in Table IV. Although ULT only uses the same number of trajectories as Oracle training, both the teacher and student policies achieve the similar performance level of Oracle performance and outperform other baseline models, which require many more trajectories to be generated for most cases. This shows the high efficiency of our proposed framework. With the increased difficulty of using omnidirectional control on all types of terrain, single-stage offline supervised transfer cannot efficiently capture the dynamics of the environment with just good trajectories, and the rough terrains make it hard for the student to survive. Although online supervised transfer and two-stage transfer significantly improve performance, ULT still outperforms supervised transfer with fewer trajectories used. Joint transfer require manual tuning of the joint ratio and struggles to handle the increased environmental challenges with similar performance to vanilla PPO with direct student-only training.
@@ VI Conclusion
Although we have simplified the training process as one single phase without training a dedicated teacher policy, we still need to retrain the model when new task requirement is raised. It is appealing to explore the continual learning and generalization capability through the power of large language models (LLMs) in handling wide variety of information with multimodality and transfer to different tasks and embodiments as future work.